### v1

# Benchmarking LLMs for Translating Classical Chinese Poetry: Evaluating Adequacy, Fluency, and Elegance

Abstract
Large language models (LLMs) have shown remarkable performance in general translation tasks. However, the increasing demand for high-quality translations that are not only adequate but also fluent and elegant. To assess the extent to which current LLMs can meet these demands, we introduce a suitable benchmark for translating classical Chinese poetry into English. This task requires not only adequacy in translating culturally and historically significant content but also a strict adherence to linguistic fluency and poetic elegance. Our study reveals that existing LLMs fall short of this task. To address these issues, we propose RAT, a Retrieval-Augmented machine Translation method that enhances the translation process by incorporating knowledge related to classical poetry. Additionally, we propose an automatic evaluation metric based on GPT-4, which better assesses translation quality in terms of adequacy, fluency, and elegance, overcoming the limitations of traditional metrics. Our dataset and code will be made available.
Benchmarking LLMs for Translating Classical Chinese Poetry:
Evaluating Adequacy, Fluency, and Elegance
Andong Chenthanks: Work was done when Andong Chen was at Pengcheng Laboratory. , Lianzhang Lou , Kehai Chenthanks:   Corresponding author. , Xuefeng Bai , Yang Xiang , Muyun Yang , Tiejun Zhao , Min Zhang School of Computer Science and Technology, Harbin Institute of Technology, China Pengcheng Laboratory, Shenzhen, China ands691119@gmail.com, {loulzh, xiangy}@pcl.ac.cn {chenkehai, baixuefeng, yangmuyun, tjzhao, zhangmin2021}@hit.edu.cn,
1 Introduction
The three difficulties in translation are: adequate, fluent, and elegant. Yan, 1898
The emergence of large language models (LLMs), especially ChatGPT, has demonstrated impressive performance in general translation tasks Tyen et al. (2023); Liang et al. (2023); Guerreiro et al. (2023); Ranaldi et al. (2023). The extensive knowledge and strong reasoning abilities of these LLMs open new possibilities in the field of translation Ouyang et al. (2022); Moslem et al. (2023b); Chen et al. (2023); He et al. (2023); Liang et al. (2023); Ghazvininejad et al. (2023); Lu et al. (2023); Chen et al. (2024a). As the demand for translation quality continues to rise, translated results must not only be adequate but also fluent and elegant Wang et al. (2024); Huang et al. (2024); Gao et al. (2024). This raises a question: can existing LLMs meet such translation requirements, and if so, to what extent can they achieve this performance?
To address this question, we introduced a suitable benchmark: translating classical Chinese poetry into English. Firstly, these poems carry culture and history, so the translation needs to adequately convey the original meaning. Secondly, classical Chinese poetry has strict rules on rhyme, tone, and structure, making fluent translation a significant challenge. Lastly, classical Chinese poetry has aesthetic value, with the concise expressions of the Chinese language showing unique elegance, which needs to be preserved in translation.
Based on the proposed classical poetry translation benchmark, current methods often lack historical and cultural knowledge, fluency, and elegance. To address these issues, we introduce RAT, a retrieval-augmented machine translation method. This method enhances the translation process by retrieving knowledge related to classical poetry, ensuring the translated output remains adequate to the original text while also achieving fluency and elegance. Furthermore, previous automatic evaluation metrics for machine translation Papineni et al. (2002); Rei et al. (2022); Sellam et al. (2020a) have primarily relied on n-gram and semantic analysis to assess entire sentences, making them unsuitable for the task of classical poetry translation. To address this issue, we propose an automatic evaluation metric based on GPT-4 (OpenAI 2024), which better assesses translation quality from the perspectives of adequacy, fluency, and elegance.
To our knowledge, this is the first evaluation and research on the translation performance of LLMs based on the task of translating classical Chinese poetry. Through this effort, we aim not only to test the capabilities of LLMs in translating classical Chinese poetry but also to inspire community discussion on the potential and future development of large models in translating texts that are adequate, fluent, and elegant.
Our contributions are summarized as follows:
  • We have introduced the first classical poetry translation dataset, which allows for better evaluation of LLMs in terms of adequacy, fluency, and elegance.
  • Based on the limitations of current LLM-based translation methods on the classical poetry translation dataset, we have proposed a retrieval-augmented translation method to further enhance the performance of large models in this task.
  • We have designed a new evaluation metric based on GPT-4 to assess classical poetry translation. This metric aligns more closely with human annotations and is better suited for evaluating the translation of classical poetry.
2 Related Work
Poetry Machine Translation The initial attempts at poetry machine translation utilized phrase-based machine translation systems to translate French poetry into metrical English poetry Genzel et al. (2010). This study demonstrated how statistical machine translation systems can produce translations while adhering to the rhythmic and rhyming rules of poetry. Chakrabarty et al. (2021) conducted an empirical study that highlighted a critical but often overlooked issue: even though they can maintain meaning and fluency, advanced machine translation systems trained on large amounts of non-poetry data struggle to preserve the original style of poetry. To address the issue of style preservation, some studies have managed to generate diverse styles by inputting sentences with specific styles into the encoder and embedding the target style into the decoder Zhang et al. (2018); Liu and Wang (2012). Considering that poetry often contains richly stylized semantic information and deep historical and cultural connotations, such as in classical Chinese poetry, Rajesh Kumar Chakrawarti and Bansal (2022) proposed a Hybrid Machine Translation (HBMT) model to improve the semantic and syntactic accuracy of poetry translations. Recently, Wang et al. (2024) have successfully implemented translations of modern poems from English to Chinese using the knowledge and multilingual capabilities of ChatGPT.
Chinese Classical Poetry Dataset Chinese classical poetry has several open datasets. Chen et al. (2019) published the first fine-grained dataset, annotating a manually fine-grained emotional poetry corpus containing 5,000 Chinese quatrains. Yutong et al. (2020) released a dataset containing 3,940 quatrains with automatically annotated themes and 1,917 quatrains with automatically annotated emotions, using a template-based method. Additionally, Liu et al. (2020) collected a parallel bilingual dataset of ancient and modern Chinese and aligned the lines using a string matching algorithm. Based on these bilingual pairs, Li et al. (2021) constructed a matching dataset to evaluate models’ semantic understanding capabilities. To our knowledge, this is the first dataset for English translation of Chinese classical poetry, aimed at evaluating the translation performance of current large models in terms of “adequate, fluent, and elegant".
3 Classical Chinese Poetry Dataset Construction
In this section, we discuss the design and construction of the MT benchmark, including the rules and steps for building this benchmark.
3.1 Discourse-Level Poetry Translation
Figure 1: An example block in the fluency and elegance in discourse-level poetry translation. The red parts indicate rhymes in both English and Chinese.
We collected a batch of classical Chinese poetry data and corresponding English translations from online resources. The classical Chinese poetry database contains 1200 poems from Tang Poems, Song Poems, and Yuan Opera. Then, We manually screened 608 classical Chinese poems and their corresponding translations to serve as gold standards for translation source and target alignment. As shown in Figure 1.
The statistics of this benchmark are shown in Table 1. We present the number of classical Chinese poetry, the number of unique tokens, the average number of tokens per sentence, and the total number of tokens in different poetry types. The source sentences in this benchmark have a moderate length, and the selected target translation sentences are well-aligned with the source in terms of length, indirectly reflecting the high quality of the reference sentences.
Poem Type Number of Poem Unique Token Average Tokens Per Sentence Total Token Numbers
Tang 197 1980/3839 11.7/13.4 11727/13115
Song 189 2214/4899 10.9/14.1 16984/18212
Yuan 222 2006/3650 12.8/13.2 12145/1197
Total 608 3059/9223 11.69/13.6 40856/42524
Table 1: Statistics on the benchmark. Numbers a/b denote the corresponding number in source sentences/target sentences.
3.2 Adequacy in Sentence-Level Translation
Figure 2: An example block in the elegance in discourse-Level poetry translation.
To construct this test set, we selected sentences containing historical knowledge and commonsense from the collected data of classical Chinese poetry. To maintain the diversity of the test set, we avoided choosing semantically similar words and those that, despite having multiple meanings in Chinese, translate to the same word in English. We preferred to select words that translate to completely different English words in different contexts, requiring disambiguation based on understanding various contexts, historical backgrounds, writing styles, and commonsense. This test set consists of 500 sentences, with each translated sentence represented as a triplet ($s$, $t_{c}$, $t_{e}$). As shown in Figure 2, $s$ contains the source sentence with ambiguous words, $t_{c}$ and $t_{e}$ are the contrast translations, with the former being correct and the latter incorrect. The selected ambiguous words are used only once in each translated sentence.
3.3 Classical Chinese Poetry Knowledge Base
Classical Chinese Poetry holds rich historical and cultural nuances, but with limited resources for Classical Chinese, modern Chinese knowledge can greatly improve translation quality. We collected a Classical Chinese Poetry Knowledge Base using open-source projects and internet resources. This Knowledge Base consists of 30,000 entries, including 30,000 Classical Chinese Poems along with their corresponding historical background, dynasty names, modern Chinese translations, author introductions, modern Chinese analysis, and poetry type information.
3.4 Human Annotation
Although collected poetry translation data were created by humans, the randomness still affects data quality. To mitigate this, we invite human annotators to assess the quality of translation parallel sentence pairs and eliminate any problematic instances.
4 Proposed Method: RAT
Figure 3: The proposed RAT framework.
Translation of classical Chinese poetry requires understanding of language style, cultural background, and author information. The RAT (Retrieval Augmented Translate) method was introduced to enhance translation outcomes by leveraging rich contextual information from the Classical Chinese Poetry Knowledge Base. We will now introduce the two workflows implemented in the RAT method in detail (in Figure 1). The first workflow (in Figure 1(a)) uses poetry as input and retrieves relevant knowledge about the poetry from the Classical Chinese Poetry Knowledge Base through text matching methods, obtaining different views of the relevant knowledge. The second workflow (in Figure 1(b)) translates based on the relevant knowledge from different views and integrates these to produce the final translation result.
4.1 The First Workflow
In the first workflow of RAT, there are two modules: Retriever and Selector. The Retriever is used to retrieve relevant knowledge from diverse perspectives based on the poetry, while the Selector filters the retrieved knowledge to enhance its relevance to the poetry.
1. Retriever. We propose a retrieval augmentation method to obtain knowledge relevant to translating classical Chinese poetry. Based on the Classical Chinese Poetry Knowledge Base we constructed, we use text-matching methods to retrieve uniquely relevant knowledge from multiple perspectives. These perspectives include historical background, dynasty name, modern Chinese translation, author introduction, modern Chinese analysis, and poetry type. The retrieved examples are shown in Appendix A.
2. Selector. The goal of the selector is to filter out irrelevant content from the results of the retriever that are unrelated to the translated sentences and passages. As an agent set by the LLM, the selector first understands the historical background, author introduction, and modern Chinese analysis based on the source sentences, and then outputs content relevant to the translated source sentences, with a limit on the output length (max_len). Specific prompts are displayed in Appendix A.1.
4.2 The Second Workflow
In the second workflow of RAT, there are three modules: Translator, Voter, and Extractor. The Translator module translates different English verses based on classical Chinese poetry and the retrieved knowledge. The Voter module then votes on each translated sentence to determine the optimal translation. Finally, the Extractor module extracts the voted translation results to form the final translated version.
1.Translator. The goal of the Translator is to translate different types of retrieved content. The LLM uses its translation capabilities to process both source content and retrieved content separately. This time, six types of related content were retrieved for classical Chinese poetry, resulting in six different translation outputs. Specific prompts are displayed in Appendix B.1.
2.Voter. The purpose of the Voter is to integrate translations based on different retrieval results to improve translation quality. As an agent of the LLM, Voter votes on each sentence of all translation outputs based on the source input, selects the highest-quality translations, and concatenates these selected sentences to form the final translation result. Specific prompts are displayed in Appendix B.2.
3.Extractor. The goal of the Extractor is to extract the final translation results generated by the Voter. As an agent of the LLM, the Extractor filters out noise from the content generated by the Voter based on the source input, and outputs the final translation results. Specific prompts are displayed in Appendix B.3.
5 Evaluation Methods
5.1 Evaluation Criteria
The translation of classical poetry requires not only artistic expression but also an understanding of the cultural background, making it nearly impossible to produce a single, universally correct translation. Therefore, multiple translations of the same classical poem can coexist, each offering a unique perspective and interpretation. Following this line of thought, we evaluate the quality of classical poetry translations through three aspects: “Adequate," “Fluent," and “Elegant." The specific evaluation methods are as follows:
Figure 4: Examples of different evaluation metrics.
  1. 1.
    Adequate
    1. (a)
      Accuracy (Acc)↑: Focus on the precision of each element in the translation, accurately translating historical, cultural, and factual aspects, including words and phrases, to maintain the correct semantic and logical relationships of the poem.
  2. 2.
    Fluent
    1. (a)
      Beauty of Sound (BS)↑: The beauty of sound in Chinese classical poetry is primarily reflected in its rhyme and meter. This standard examines whether the translation achieves harmonious sound, precise wording, adherence to strict metrical rules, and a rhythm that is both smooth and dynamic. As shown in Figure 4(b).
    2. (b)
      Beauty of Form (BF)↑: Chinese classical poetry emphasizes balanced structure, with common forms such as "four-character quatrains," "seven-character quatrains," and "seven-character regulated verse," each exhibiting distinct aesthetic characteristics of Chinese poetry. This standard evaluates whether the translation maintains consistency with the source poem’s structure, including the alignment of line numbers and balanced phrasing. As shown in Figure 4(c).
  3. 3.
    Elegant
    1. (a)
      Beauty of Meaning (BM)↑: Chinese classical poetry uses concise and precise language to create vivid imagery and a rich atmosphere for readers. This standard assesses the depth and richness of the translation, focusing on the effectiveness of conveying themes, emotions, and messages. As shown in Figure 4(a).
5.2 LLM-based Classical Poetry Metric
In this section, we aim to establish a method for evaluating the translation of classical Chinese poetry based on LLMs assessments. Drawing from current QE (Quality Estimation) research, we have developed an approach to utilize LLM in determining the quality of classical Chinese poetry translations. As shown in Figure 5, we designed a 1-5 scoring prompt to help LLM focus on the translation quality in terms of Beauty of Sound (LLM-BS). In the evaluation criteria, a score of 1 reflects a poor translation of classical Chinese poetry, a score of 3 indicates a basic but imperfect translation result and a score of 5 represents an excellent and accurate translation in this aspect. Subsequently, test data is provided to the LLM, and the LLM is asked to generate an assessment with scores only. Additionally, we calculate the average scores of BF, BF, and BM (LLM-Avg) to evaluate the overall translation performance. The prompt details for Beauty of Form (LLM-BF) and Beauty of Meaning (LLM-BM) are provided in Appendix F and G.
{mdframed}
[backgroundcolor=chinese_red, linecolor=chinese_red, linewidth=2pt, roundcorner=10pt] /* Task prompt */
Evaluate the beauty of sound in the given Chinese translation of classical poetry. Focus on whether the translation achieves harmonious sound, precise wording, adherence to strict metrical rules, and a rhythm
1 point: Poor translation, lacks harmony and adherence to metrical rules, and fails to capture the beauty of sound.
2 point: Below average, some rhyme and meter present but with noticeable imperfections and awkwardness.
3 point: Basic translation, captures some aspects of sound beauty but with several imperfections in rhyme, meter, or rhythm.
4 point: Good translation, mostly harmonious with minor imperfections in sound quality or adherence to metrical rules.
5 point: Excellent translation, achieves harmonious sound, precise wording, strict adherence to metrical rules, and a smooth, dynamic rhythm.
/* Input Data */: {mdframed}[backgroundcolor=chinese_red_small, linecolor=chinese_red_small, linewidth=2pt, roundcorner=10pt] Original Chinese poem: {source}
English translation: {translation}
Evaluation (score only):
/*Output Text */: {mdframed}[backgroundcolor=chinese_red_small, linecolor=chinese_red_small, linewidth=2pt, roundcorner=10pt] {score}
Figure 5: Evaluation of the beauty of sound in Chinese translation of classical poetry
5.3 Traditional Automatic Evaluation
In evaluating our translation methodology, we initially employ COMET [^1] https://huggingface.co/Unbabel/wmt22-comet-da Rei et al. (2022) and BLEURT [^2] https://github.com/lucadiliello/bleurt-pytorch Sellam et al. (2020b) as automatic metrics, aligning with the established standards in LLM-based translation literature He et al. (2023); Huang et al. (2024). For traditional translation evaluation, we use BLEU  [^3] https://github.com/mjpost/sacrebleu Papineni et al. (2002).
6 Experiment Setup
6.1 Comparing Systems
In our evaluation, RAT is compared with a range of translation methods, including Zero-shot Wei et al. (2022), 5-shot Brown et al. (2020); Hendy et al. (2023), Rerank Moslem et al. (2023a), Refine Chen et al. (2023), MAD Liang et al. (2023), and MAPS He et al. (2023). To validate its generalizability, we utilize three LLMs, which include closed-source models such as ChatGPTOuyang et al. (2022) and GPT-4Achiam et al. (2023)  [^4] The ChatGPT and GPT-4 models used in this work are accessed through the gpt-3.5-turbo and gpt-4 APIs, respectively., as well as open-source models like Alpaca-7B Taori et al. (2023)  [^5] https://huggingface.co/tatsu-lab/alpaca-7b-wdiff/tree/main and Vicuna-7B Chiang et al. (2023)  [^6] https://huggingface.co/lmsys/vicuna-7b-v1.5. The following content will provide detailed descriptions of these comparisons:
  • Baseline, standard zero-shot translation is performed in ChatGPT Ouyang et al. (2022) and GPT-4 Achiam et al. (2023). The temperature parameter set to 0, which is the default value for our experiments.
  • 5-Shot Hendy et al. (2023), involves prepending five high-quality labelled examples from the training data to the test input.
  • Rerank Moslem et al. (2023a) was conducted with the identical prompt as the baseline, employing a temperature of 0.3 (Moslem et al., 2023b). Three random samples were generated and combined with the baseline to yield four candidates. The optimal candidate was chosen through GPT-4.
  • Renfie Chen et al. (2023) first requests a translation from ChatGPT, then provides the source text and translation results, and obtains a refined translation through multiple rounds of modifications by mimicking the human correction process.
  • MAD Liang et al. (2023) enhance the capabilities of large language models (LLMs) by encouraging divergent thinking. In this method, multiple agents engage in a debate, while an agent oversees the process to derive a final solution.
  • Dual-ReflectChen et al. (2024b) provide supervisory signals for large models to reflect on translation results through dual learning, thereby iteratively improving translation performance (the maximum number of iterations is set to 5).
  • RAT,is proposed method in this work.
7 Experimental Results
7.1 Can LLM evaluate Chinese classical Poetry ?
We first randomly select N chapter-level Chinese classical poetry pieces from each source language dataset and translate them using the RAT method. We then compare the evaluation results of the large language model with human-annotated results to calculate the Pearson correlation coefficient Pearson (1920), Spearman correlation coefficient Spearman (1961), and Kendall correlation coefficient Kendall (1948) to determine the level of consistency.
Metric Pearson’s $r$ Spearman’s $\rho$ Kendall’s $\tau$
BLEU -0.23 -0.18 -0.12
BLEU-1 0.05 0.08 0.05
BLEURT 0.14 0.16 0.11
COMET 0.13 0.18 0.11
LLM-BM 0.85 0.81 0.85
LLM-BF 0.71 0.75 0.70
LLM-BS 0.73 0.73 0.76
LLM-AVG 0.77 0.73 0.75
The results in Table N indicate that large language models can serve as effective tools for evaluating the translation quality of idioms across different language pairs. In contrast, BLEU, COMET, and BLEURT were found to be less consistent with human evaluations, highlighting the advantages of our proposed evaluation method in the translation of classical poetry.
7.2 Main Results
Methods Discourse-Level Poetry Transaltion
COMET BLEURT LLM-BM LLM-BS LLM-BF LLM-Avg BLEU BLEU-1 BLEU-2 BLEU-3 BLEU-4
GPT-4 60.3 43.0 4.0 3.7 3.6 3.8 4.0 22.1 7.8 3.3 1.7
ChatGPT 61.1 42.4 3.3 3.2 2.9 3.1 4.3 23.4 8.7 3.1 1.8
   +5shot 61.0 42.5 3.5 3.3 3.3 3.4 3.7 22.0 7.7 3.2 1.6
   +rerank 61.0 42.5 3.7 3.7 3.9 3.8 3.7 22.5 8.0 3.4 1.7
   +MAD 59.9 42.3 3.7 3.6 3.8 3.8 4.3 23.2 8.8 3.7 1.8
   +Dual-Reflect 58.2 40.9 3.8 3.8 3.9 3.9 3.8 20.5 7.5 3.2 1.6
   +EAPMT 61.1 42.9 3.8 3.7 3.8 3.7 3.7 21.6 7.5 3.1 1.5
   +RAT 62.7 43.9 4.1 3.9 3.9 4.0 4.6 23.9 9.8 3.9 2.2
Llama3-8B 54.3 37.4 2.7 2.6 2.4 2.5 1.7 17.4 6.1 3.5 1.3
   +5shot 54.5 37.6 2.9 2.8 2.6 2.7 1.8 17.4 6.2 3.4 1.3
   +Rerank 54.8 38.1 3.0 3.3 3.5 3.2 2.0 17.9 6.6 3.6 1.5
   +RAT 55.6 38.4 3.4 3.3 3.6 3.4 2.2 18.2 7.0 3.9 1.8
Vicuna-7B 52.2 26.4 2.4 2.4 1.8 2.2 0.3 16.5 4.7 3.4 1.0
   +5shot 52.4 26.1 2.5 2.6 2.3 3.6 1.0 17.1 4.3 3.6 1.3
   +Rerank 52.8 26.3 3.0 2.6 3.3 2.8 1.2 17.5 5.0 3.7 1.6
   +RAT 60.1 26.9 3.0 2.5 3.3 2.9 1.8 17.6 5.3 3.9 1.9
We compare various different LLM-based methods on the Classical Chinese Poetry Benchmark with RAT. The analysis of the experimental results is as follows:
  • The task of translating Classical Chinese Poetry is challenging. The experiment shows that translating Classical Chinese Poetry is difficult, with low COMET/BLEURT/BLEU scores for current LLM-MT methods. This is because the translation is challenging and n-gram-based evaluation methods are not ideal for poetry. Additionally, GPT-4-based evaluation metrics, particularly in BS/BM/BF/AVG aspects, still need considerable improvement.
  • The effectiveness of RAT method. The proposed RAT method consistently outperforms other comparative methods across all evaluation metrics, demonstrating the effectiveness of our approach.
  • The effectiveness of different LLMs. Among all comparative methods, closed-source models perform better on this task than open-source models, possibly implying that closed-source models benefit from richer pre-training data, thus enabling higher-quality translations. This also suggests that the Chinese Classical Poetry translation task is more challenging.
  • The effectiveness of retrieved knowledge. For EAPMT, which relies on LLMs’ self-generated knowledge, the RAT method, which uses retrieval-based knowledge, provides more accurate information and further improves translation quality. This suggests that translating Chinese Classical Poetry demands stricter knowledge. However, the results from RAT still do not reach the level of human translation, indicating significant room for improvement in this task.
7.3 Evaluation of Adequacy
To evaluate the translation performance of LLMs in terms of Adequacy, we employed a constructed dataset of 500 Classical Chinese Sentence-Level Translations to assess various translation methods. This experiment follows the method of Liang et al., 2023, assessing translation results from two main dimensions: manual evaluation of ambiguous translation accuracy (see Appendix A for details) and direct evaluation of translation quality using the LLM-BM.
The experimental results are shown in Table 2. In terms of ambiguity resolution accuracy, RAT performed the best, indicating that retrieving accurate information helps to better eliminate ambiguities in translation tasks. In the human evaluation, RAT received the highest score, further demonstrating that this method achieves excellent translation quality.
Methods LLM-BM ACC
GPT-4 3.5 57.4
ChatGPT
   +Zero-Shot 2.9 51.8
   +Rerank 3.0 56.4
   +Dual-Reflect 3.3 54.3
   +MAD 3.5 56.1
   +RAT 3.8 57.8
Table 2: The human-annotated results of the Adequacy in Sentence-Level Commonsense MT benchmark Translation.
8 Analysis
8.1 LLM-based Metric Consistency
In this experiment, we evaluated whether the proposed LLM-based metrics, including LLM-BS, LLM-B, LLM-BM, and LLM-AVG, could reflect the evaluation characteristics of Beauty of Sound, Beauty of Form, Beauty of Meaning, and overall translation quality, while also demonstrating variability. We conducted pairwise correlation tests between human evaluation results and LLM-based results, employing the Pearson correlation coefficient, Spearman correlation coefficient, and Kendall correlation coefficient. The experimental results are illustrated in Figure N.
Figure 6: LLM-based Metric Consistency Experiment. In the heatmap, the horizontal axis represents the human evaluation results, and the vertical axis represents the LLM evaluation results.
The experimental results indicate that, among all correlation coefficients, the consistency results based on the same annotations are significantly higher than the other results. This demonstrates the rationality of the evaluation settings for LLM-BS, LLM-B, LLM-BM, and LLM-AVG in the experiment.
8.2 Impact of Different Knowledge on Translation Performance
The RAT method utilized knowledge from the Classical Chinese Poetry Knowledge Base for translation. To investigate which specific knowledge is more effective in translating Classical Chinese Poems within knowledge-based translation methods, this experiment employed a modified RAT method that used only one type of retrieved knowledge at a time and removed the Voter module (Figure 4). The experimental results are presented in Figure 7.
Figure 7: Experiment on the Impact of Different Knowledge of Classical Chinese Poetry on Translation
On the one hand, retrieval-based translation methods showed overall performance improvement, indicating that knowledge plays a significant role in the task of translating Classical Chinese Poetry. On the other hand, the results revealed that knowledge of modern Chinese translation contributed the most to performance enhancement. This suggests that there are differences in form, semantics, and rhythm between Classical Chinese Poems and modern Chinese, further demonstrating the considerable challenges of directly translating Classical Chinese Poetry.
8.3 Translation Challenges Across Different Types of Classical Chinese Poetry
Figure 8: Experiment on the Impact of Different Types of Classical Chinese Poetry on Translation
To explore the translation difficulty of different types of poems (Tang, Song, and Yuan) in the collected 608 Chinese Classical Poems, we employed the RAT method to generate translation results for each poetry type. These results were then evaluated using LLM-BF, LLM-BM, LLM-BS, and LLM-AVG. The experimental results are shown in Figure 8.
From the experimental results, we can observe that the translation trends are consistent across different types. Firstly, the translation of Tang poetry is relatively easier compared to other types, likely due to its stricter format and fewer words. Secondly, the lower scores of LLM-BF and LLM-BS indicate that the translation of poetic structure and rhythm remains a significant challenge for LLMs. Lastly, the generally higher results of LLM-BM suggest that retrieval-based methods can produce more adequate translations.
9 Conclusion
Our research reveals the challenges LLMs face in translating classical Chinese poetry, especially in cultural knowledge, fluency, and elegance. We introduced the Retrieval-Augmented Translation (RAT) method and a GPT-4-based evaluation metric to improve translation quality. This first comprehensive evaluation exposes LLMs’ limitations and suggests areas for improvement, aiming to inspire better culturally informed translations.
10 Acknowledgements
Appendix A Human Evaluations
In this section, we conduct a human evaluation to measure translation quality. We assess ambiguity resolution. Four native English speakers were invited to participate. In the Adequacy sentence-level task, the four experts scored each sentence for ambiguity resolution against the reference, awarding 1 point for resolved and 0 points for unresolved.
A.1 Detailed prompt for Selector
{mdframed}
[backgroundcolor=purple!10, linecolor=white, linewidth=2pt, roundcorner=10pt]Part-1: Selector: Please identify the knowledge related to {max_len}-word content to translating this classical Chinese poem {text} from the {rag context} knowledge base.
Input Text: {mdframed}[backgroundcolor=blue!10, linecolor=purple!10, linewidth=2pt, roundcorner=10pt]Source Poem, Sentence Length and Retrieved knowledge
Output Text: {mdframed}[backgroundcolor=yellow!10, linecolor=purple!10, linewidth=2pt, roundcorner=10pt] Refined knowledge.
Appendix B Detail Prompt
B.1 Detailed prompt for Translator
{mdframed}
[backgroundcolor=purple!10, linecolor=white, linewidth=2pt, roundcorner=10pt]Part-2: Translator: Please translate this classial a Chinese poem {translate type} into a English poem {translate type}: Explanation:{rag context} Poem:{text}
Input Text: {mdframed}[backgroundcolor=blue!10, linecolor=purple!10, linewidth=2pt, roundcorner=10pt]Source Poem, Retrieved knowledge and Potery Type
Output Text: {mdframed}[backgroundcolor=yellow!10, linecolor=purple!10, linewidth=2pt, roundcorner=10pt] Translated English Poem
B.2 Detailed prompt for Voter
{mdframed}
[backgroundcolor=purple!10, linecolor=white, linewidth=2pt, roundcorner=10pt]Part-3: Iterative Refinement: Using the classical Chinese poem {src_text} as a source, compare six translation candidates to determine the highest quality result. Avoid including unrelated content. Here are the candidates: First, {s1}; second, {s2}; third, {s3}; fourth, {s4}; fifth, {s5}; sixth, {s6}.
Input Text: {mdframed}[backgroundcolor=blue!10, linecolor=purple!10, linewidth=2pt, roundcorner=10pt]Source Sentence, Translated Resluts based on six knowledge
Output Text: {mdframed}[backgroundcolor=yellow!10, linecolor=purple!10, linewidth=2pt, roundcorner=10pt] Translated Result
B.3 Detailed prompt for Extractor
{mdframed}
[backgroundcolor=purple!10, linecolor=white, linewidth=2pt, roundcorner=10pt]Part-4:Understanding-Based Translation: Extract only translation-relevant content from {target text} based on {text}. Input Text: {mdframed}[backgroundcolor=blue!10, linecolor=purple!10, linewidth=2pt, roundcorner=10pt] The final translation result.
Output Text: {mdframed}[backgroundcolor=yellow!10, linecolor=purple!10, linewidth=2pt, roundcorner=10pt]Target Sentence $t$
B.4 Detailed prompt for Beauty of Meaning
{mdframed}
[backgroundcolor=chinese_red, linecolor=chinese_red, linewidth=2pt, roundcorner=10pt] /* Task prompt */
Evaluate the translation of Chinese classical poetry for the beauty of meaning, focusing on whether the translation effectively conveys the themes, emotions, and messages of the original. This includes the use of concise and precise language to create vivid imagery and a rich atmosphere.
1 point: Poor translation, fails to convey the depth and richness of the original poetry.
2 point: Basic translation with significant shortcomings in capturing themes, emotions, and messages.
3 point: Satisfactory translation, conveys basic themes and emotions but lacks refinement or depth.
4 point: Good translation, effectively captures most themes, emotions, and messages with minor imperfections.
5 point: Excellent translation, accurately conveys the depth, richness, and atmosphere of the original poetry with full thematic and emotional resonance.
/* Input Data */: {mdframed}[backgroundcolor=chinese_red_small, linecolor=chinese_red_small, linewidth=2pt, roundcorner=10pt] Original Chinese poem: {source}
English translation: {translation}
Evaluation (score only):
/*Output Text */: {mdframed}[backgroundcolor=chinese_red_small, linecolor=chinese_red_small, linewidth=2pt, roundcorner=10pt] {score}
Figure 9: Evaluation of the beauty of meaning in Chinese translation of classical poetry
B.5 Detailed prompt for Beauty of Form
{mdframed}
[backgroundcolor=chinese_red, linecolor=chinese_red, linewidth=2pt, roundcorner=10pt] /* Task prompt */
Evaluate the translation of the given Chinese classical poem into English. Focus on whether the translation maintains consistency with the source poem’s structure, including the alignment of line numbers and balanced phrasing.
1 point: Poor translation, disregards the poem’s structure, and fails to convey its aesthetic qualities.
2 point: Some attempt to maintain structure but lacks alignment and aesthetic consistency.
3 point: Basic structural elements are maintained but with noticeable imperfections in alignment and phrasing.
4 point: Good translation, with most structural elements preserved and minor issues in phrasing and alignment.
5 point: Excellent translation, accurately preserving the structure, alignment, and aesthetic qualities of the original poem.
/* Input Data */: {mdframed}[backgroundcolor=chinese_red_small, linecolor=chinese_red_small, linewidth=2pt, roundcorner=10pt] Original Chinese poem: {source}
English translation: {translation}
Evaluation (score only):
/*Output Text */: {mdframed}[backgroundcolor=chinese_red_small, linecolor=chinese_red_small, linewidth=2pt, roundcorner=10pt] {score}
Figure 10: Evaluation of the beauty of form in Chinese translation of classical poetry